# Batched Bayesian optimization with correlated candidate uncertainties

### v1

@@ Abstract
Batched Bayesian optimization (BO) can accelerate molecular design by efficiently identifying top-performing compounds from a large chemical library. Existing acquisition strategies for batch design in BO aim to balance exploration and exploitation. This often involves optimizing non-additive batch acquisition functions, necessitating approximation via myopic construction and/or diversity heuristics. In this work, we propose an acquisition strategy for discrete optimization that is motivated by pure exploitation, qPO (multipoint Probability of Optimality). qPO maximizes the probability that the batch includes the true optimum, which is expressible as the sum over individual acquisition scores and thereby circumvents the combinatorial challenge of optimizing a batch acquisition function. We differentiate the proposed strategy from parallel Thompson sampling and discuss how it implicitly captures diversity. Finally, we apply our method to the model-guided exploration of large chemical libraries and provide empirical evidence that it performs better than or on par with state-of-the-art methods in batched Bayesian optimization.
@@ 1 Introduction
Predictive modeling can greatly accelerate molecular and materials discovery. Data-driven and simulation-based models can aid in prioritizing experiments for the design of drugs (Liu et al., 2024; 2023; Horne et al., 2024), light-emitting diodes (Gómez-Bombarelli et al., 2016), and other materials (Schwalbe-Koda et al., 2021). In iterative design cycles guided by predictive models, the reliability of model predictions can be gradually improved as the model is updated with newly collected data. This iterative, model-guided approach has been applied to the discovery of drugs (Desai et al., 2013), laser emitters (Strieth-Kalthoff et al., 2024), and dyes (Koscher et al., 2023).
@@
Bayesian optimization (BO) is arguably the most popular mathematical framework for iterative model-based design (Frazier, 2018; Garnett, 2023). BO optimizes an expensive black-box objective function ($f$ or “oracle”) by iteratively (re)training a surrogate model and using its predictions to select designs for evaluation (Frazier, 2018). At each iteration, an acquisition function uses the mean and/or uncertainty of surrogate model predictions to select which design(s) to evaluate. BO has been applied to efficiently explore chemical libraries in numerous previous works (Cherkasov et al., 2006; Graff et al., 2021; Yang et al., 2021; Bellamy et al., 2022; Wang-Henderson et al., 2023).
@@
Computational (e.g., physics-based simulations) and experimental (e.g., bioactivity assays) oracles in many chemistry applications are most efficiently evaluated in parallel. The total evaluation budget is therefore spread across relatively few iterations with large batch sizes, requiring an acquisition function to select a $batch$ of experiments in each iteration. The non-additivity of batch-level acquisition functions complicates the selection of optimal batches in BO; when the value of selecting a candidate depends on other selections, batch design becomes a combinatorial problem. Optimizing Bayes-optimal batch acquisition functions is therefore often computationally intractable even for modest batch sizes (Garnett, 2023, Section 11.3). Strategies that fail to consider this non-additivity, such as selecting the top candidates based on a sequential policy, can produce homogeneous batches that lack diversity. Prior works have primarily relied on (1) heuristics to increase diversity (Gonzalez et al., 2016; Kathuria et al., 2016; Nava et al., 2022; Nguyen et al., 2016; Groves & Pyzer-Knapp, 2018), (2) hallucinated observations to approximate intractable integrals (Ginsbourger et al., 2010; Desautels et al., 2014), or (3) the randomness inherent to Thompson sampling (Thompson, 1933) to extend it to the batch setting (Hernández-Lobato et al., 2017; Dai et al., 2022; Ren & Li, 2024).
@@
Most batch acquisition strategies aim to optimally balance exploration and exploitation in the limit of infinite iterations. In settings with few iterations, there are fewer opportunities to refine designs found from exploration, and prioritizing exploitation may be desirable. While exploration is expected to prevent the selection of homogeneous batches, we hypothesized that diversity is reflected by the surrogate model posterior (e.g., as a lack of covariance between predictions) and can be achieved without reliance on exploration. In this paper, we propose qPO (multipoint Probability of Optimality), a batch construction strategy that maximizes the likelihood that the optimum exists in the acquired batch. Inspired by parallel Thompson sampling (Hernández-Lobato et al., 2017), qPO centers around the probability of optimality, accounts for correlations between inputs, and is naturally parallelizable. However, qPO aims to forego randomness in the pursuit of pure exploitation. Uniquely, the defined batch-level acquisition function can be expressed as a sum of individual candidate acquisition scores, circumventing the combinatorial challenge of maximizing a batch-level acquisition function. We summarize the contributions of this work as follows:
@@
    We demonstrate that our acquisition strategy identifies top-performers from chemical libraries as or more efficiently than state-of-the-art alternatives for batched BO in two realistic molecular discovery settings.
@@ 2.1 Preliminaries
We first assume that there exists an expensive black box oracle function $f(\cdot)$ that maps each candidate $x_{i}$ to a scalar objective value $y_{i}$. Following Hernández-Lobato et al. (2017), we assume that evaluations of $f$ are noise-free. Our aim is to solve the following optimization problem:
@@
where $\mathcal{X}$ is a fixed discrete design space comprised of $N$ candidates $\{x_{i}\}_{i=1,...,N}$.
Bayesian optimization (BO) aims to solve this optimization using an iterative model-guided approach. In each iteration, we select $b$ candidates for parallel evaluation with the oracle $f$. We denote the set of acquired candidates $\mathcal{X}_{acq}$. In each iteration, $f$ is evaluated for all $x_{i}$ in $\mathcal{X}_{acq}$, and a surrogate model $\hat{f}$ that predicts $f$ is (re)trained with newly acquired data. An acquisition function utilizes surrogate model predictions on $x_{i}$ for $i=1,...,N$ to select the next set of evaluations. The iterative procedure ends when some stopping criterion is met: computational or experimental resources are expended, a maximum number of iterations is reached, or a satisfactory value of $f$ is achieved.
The imperfect surrogate model $\hat{f}$ provides a probabilistic prediction of the objective function value(s) for one or more candidates. This distribution over possible values of $y$ for candidate $x$, denoted $p(y|x)$, may be described by a machine learning model such as a Gaussian process or Bayesian neural network. We may alternately have a surrogate model that does not form a continuous probability distribution but still enables sampling, e.g., through deep ensembling (Lakshminarayanan et al., 2017) or Monte Carlo dropout (Gal & Ghahramani, 2016). Without loss of generality, we consider the prediction to be an integral over possible model parameters $\bm{\theta}$:
@@
where $p(\bm{\theta})$ does not truly have to represent a prior, but can represent a posterior distribution given some training data. We consider all candidates $x_{i}\in\mathcal{X}$ to be deterministic and fixed.
Figure 1: Exploitative batch design by maximizing the likelihood of including the optimum, in the context of a single Bayesian optimization iteration. First, a surrogate model (or ensemble) is trained on acquired data. Second, samples are obtained from the joint posterior distribution over all candidates. When direct posterior sampling is impossible or inefficient, a multivariate Gaussian may be modeled from the true posterior to enable approximate posterior sampling. Third, we estimate from these samples the probability that each candidate is the true optimum. Fourth, the batch is populated with candidates most likely to be optimal; in doing so, the proposed strategy maximizes the probability that the batch contains the true optimum. In addition to the sampling-based approach visualized here, we describe alternative methods to approximate acquisition scores in Section 2.3.2.
@@ 2.2 Deriving an acquisition function for optimal batch design
Most acquisition strategies in BO aim to balance exploitation and exploration. Exploitation prioritizes selections that are most likely to achieve the highest oracle score, while exploration is intended to prevent a search from getting “stuck” in local optima. Conceptually, exploration is expected to contribute to a more reliable surrogate model and thereby benefit the optimization in the long run; a classic failure mode of BO within continuous design spaces is the oversampling of a single mode of the surrogate model posterior (Hernández-Lobato et al., 2014). In contrast, exploitation selects the best candidates at a given iteration without consideration of the impact on future iterations. We pursue a batch acquisition strategy that is purely exploitative, optimizing expected performance in the immediate iteration as if the optimization could be stopped at any time. [^1] This is a realistic setting for molecular discovery campaigns. When the cost of evaluating $f$ varies across compounds in the design space, the number of iterations or oracle budget may be uncertain when the optimization begins. Our exploitative acquisition strategy is visualized within the context of one BO iteration in Figure 1.
@@
We refer to this acquisition function as qPO (multipoint Probability of Optimality), which can be equivalently defined as:
$\Pr(x^{*}\in\mathcal{X}_{acq})=\sum_{x_{i}\in\mathcal{X}_{acq}}\Pr(x^{*}=x_{i}\,|\,\{x_{j}\neq x^{*}\}_{x_{j}\in\mathcal{X}_{acq},j\neq i}),$ (4)
where the probability of each selection being optimal is conditioned on all others not being optimal. We focus on continuous surrogate models with low but non-zero observation noise (e.g., Gaussian processes), such that the probability of two inputs having the exact same output is 0. Therefore, we may assume that the events $\{\hat{f}(x^{*})=\hat{f}(x_{i})\}_{x_{i}\in\mathcal{X}}$ are mutually exclusive. In our estimation of Eq. 4, we will treat $\{\hat{f}(x^{*})=\hat{f}(x_{i})\}$ as equivalent to $\{x^{*}=x_{i}\}$, leading to the presumed mutual exclusivity of the events $\{x^{*}=x_{i}\}_{x_{i}\in\mathcal{X}}$ and the following result:
$\Pr(x^{*}\in\mathcal{X}_{acq})=\sum_{x_{i}\in\mathcal{X}_{acq}}\Pr(x^{*}=x_{i}).$ (5)
We elaborate on scenarios where a unique optimum cannot be assumed in Appendix A.1. An appropriate strategy to optimize the objective in Eq. 5 is to approximate $\Pr(x^{*}=x_{i})$ for each candidate $x_{i}$ and select the top $b$ candidates based on the resulting approximations.
This acquisition score is deceptively simple to apply in practice. For a given $x_{i}$ in a list of candidates $\mathbf{x}$, we may evaluate $\Pr(x^{*}=x_{i})$ as the following equivalent integral or expectation:
$\Pr(x^{*}=x_{i})$ $=P(f(x^{*})=f(x_{i}))$ (6)
$=P(y_{i}>y_{j}~{}\forall~{}j\neq i)$ (7)
$=P(y_{i}=\text{max}(\mathbf{y})\,)$ (8)
@@
$=\mathbb{E}_{\mathbf{y}\sim p(\mathbf{y}|\mathbf{x})}\left[\mathbbm{1}_{y_{i}=\text{max}(\mathbf{y})}\right].$ (10)
A special outcome of this batch design strategy is that qPO is naturally a sum over individual acquisition scores. Batch acquisition functions that are extensions of sequential ones, such as multipoint expected improvement (qEI) (Ginsbourger et al., 2010), typically require evaluation of the entire batch at once. Myopic construction of large batches can only approximately optimize these batch-level acquisition functions. In contrast, due to the mutual exclusivity of the events $\{x^{*}=x_{i}\}$ under the circumstances defined previously, qPO can be optimized as a batch-level acquisition function without the issue of non-additivity and resulting approximations faced by many myopic batch construction strategies. Additionally, a sequential qPO policy would be identical to the batched one, acquiring the single candidate with the greatest probability of being the optimum.
@@ 2.3.1 Monte Carlo integration to approximate acquisition scores
The expectation in Eq. 10 may be approximated with $M$ samples from the surrogate model posterior:
$\Pr(x^{*}=x_{i})$ $\approx\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}_{y_{i}^{(m)}=\text{max}(\mathbf{y}^{(m)})}.$ (11)
In practice, these may be samples from the posterior of a Gaussian process, predictions based on sampled parameters $\bm{\theta}^{(m)}$ from a Bayesian neural network, or predictions based on Monte Carlo dropout. The result in A.2 shows that candidates with a high probability of optimality are very likely to appear in the batch even with modest $M$, making a Monte Carlo estimate reasonable. Algorithm 1 summarizes the implementation of qPO using Monte Carlo sampling.
@@
     Compute joint posterior $p(\mathbf{y}|\mathbf{x},\bm{\theta})$ over candidates $\bf{x}$
@@
     Evaluate $y_{i}=f(x_{i})$ for all $x_{i}\in\mathcal{X}_{acq}^{*}$triangleright Call the oracle
@@
The only requirement for this approximation of acquisition scores is that the posterior can be sampled. However, efficient and widespread algorithms exist to sample from Gaussian distributions in particular (Vono et al., 2022; Aune et al., 2013). In practice, the computational cost of posterior sampling may be reduced by sampling from a multivariate Gaussian that approximates the true posterior. This approximate posterior may be obtained from an arbitrary number of samples from the true posterior. For example, one might obtain a multivariate Gaussian posterior based on the predictions of a neural network ensemble; this could apply for an ensemble of any model architecture and for an arbitrary number of models in the ensemble. Specifically, given a smaller number ($<M$) of samples, one can construct an empirical mean and covariance matrix from which additional samples can be drawn.
@@
If few candidates are perceived by the model as having a substantial probability of being optimal, $\Pr(x^{*}=x_{i})$ may be estimated to be zero for many candidates for finite $M$. Filling a batch of size $b$ may emerge as a challenge if there are fewer than $b$ non-zero acquisition scores. A simple way to address this issue is to fill the remainder of the batch using an alternative metric (e.g., greedy or upper confidence bound). Methods designed for rare event estimation (de Boer et al., 2005; Cérou et al., 2012; Gibson & Kroese, 2022) may alternatively be implemented to assign acquisition scores to candidates with small probabilities of optimality.

### v2

# Batched Bayesian optimization by maximizing the probability of including the optimum

@@ Abstract
Batched Bayesian optimization (BO) can accelerate molecular design by efficiently identifying top-performing compounds from a large chemical library. Existing acquisition strategies for batch design in BO aim to balance exploration and exploitation. This often involves optimizing non-additive batch acquisition functions, necessitating approximation via myopic construction and/or diversity heuristics. In this work, we propose an acquisition strategy for discrete optimization that is motivated by pure exploitation, qPO (multipoint Probability of Optimality). qPO maximizes the probability that the batch includes the true optimum, which is expressible as the sum over individual acquisition scores and thereby circumvents the combinatorial challenge of optimizing a batch acquisition function. We differentiate the proposed strategy from parallel Thompson sampling and discuss how it implicitly captures diversity. Finally, we apply our method to the model-guided exploration of large chemical libraries and provide empirical evidence that it is competitive with and complements other state-of-the-art methods in batched Bayesian optimization.
[…]
@@ 1 Introduction
Predictive modeling can greatly accelerate molecular and materials discovery. Data-driven and simulation-based models can aid in prioritizing experiments for the design of drugs 1, 2, 3 and other materials 4, 5. In iterative design cycles guided by predictive models, the reliability of model predictions can be gradually improved as the model is updated with newly collected data. This iterative, model-guided approach has been applied to the discovery of drugs 6, laser emitters 7, and dyes 8, 9.
@@
Bayesian optimization (BO) is arguably the most popular mathematical framework for iterative model-based design 10, 11. BO optimizes an expensive black-box objective function ($f$ or “oracle”) by iteratively training a surrogate model and using its predictions to select designs for evaluation 10. At each iteration, an acquisition function uses the mean and/or uncertainty of surrogate model predictions to select which design(s) to evaluate. BO has been applied to efficiently explore chemical libraries in numerous previous works 12, 13, 14, 15, 16.
@@
Computational (e.g., physics-based simulations) and experimental (e.g., bioactivity assays) oracles in many chemistry applications are most efficiently evaluated in parallel. The total evaluation budget is therefore spread across relatively few iterations with large batch sizes, requiring an acquisition function to select a $batch$ of experiments. The non-additivity of batch-level acquisition functions complicates the selection of optimal batches in BO; when the value of selecting a candidate depends on other selections, batch design becomes a combinatorial problem. Optimizing Bayes-optimal batch acquisition functions is therefore often computationally intractable even for modest batch sizes 11. Strategies that fail to consider this non-additivity, such as selecting the top candidates based on a sequential policy, can produce homogeneous batches that lack diversity. Prior works have primarily relied on (1) methods to increase diversity 17, 18, 19, 20, (2) hallucinated observations to approximate intractable integrals 21, 22, (3) the randomness inherent to Thompson sampling 23 to extend it to the batch setting 24, 25, or (4) some combination thereof 26, 27.
@@
In this paper, we propose qPO (multipoint Probability of Optimality), a batch construction strategy that maximizes the likelihood that the optimum exists in the acquired batch. Inspired by parallel Thompson sampling 24, 28, qPO centers around the probability of optimality, accounts for correlations between inputs, and is naturally parallelizable. However, qPO aims to forego randomness. While this distinction may seem to hinder exploration, the consideration of a surrogate model’s joint distribution over all candidates allows qPO to favor diversity. Uniquely, the defined batch-level acquisition function can be expressed as a sum of individual candidate acquisition scores, circumventing the combinatorial challenge of maximizing a batch-level acquisition function. We summarize the contributions of this work as follows:
@@
    We demonstrate that our acquisition strategy identifies top-performers from chemical libraries as efficiently as state-of-the-art alternatives for batched BO in two realistic molecular discovery settings.
@@ 2.1 Preliminaries
We first assume that there exists an expensive black box oracle function $f(\cdot)$ that maps each candidate $x_{i}$ to a scalar objective value $y_{i}$. Following Hernández-Lobato et al. 24, we assume that evaluations of $f$ are noise-free. Our aim is to solve the following optimization problem:
@@
where $\mathcal{X}$ is a fixed discrete design space, e.g., of molecular structures from a virtual library comprised of $N$ candidates $\{x_{i}\}$ for $i=1,...,N$.
Bayesian optimization (BO) aims to solve this optimization using an iterative model-guided approach. In each iteration, we select $b$ candidates for parallel evaluation with $f$. We denote the set of acquired candidates $\mathcal{X}_{acq}$. In each iteration, $f$ is evaluated for all $x_{i}$ in $\mathcal{X}_{acq}$, and a surrogate model $\hat{f}$ that predicts $f$ is (re)trained with newly acquired data. An acquisition function utilizes surrogate model predictions on candidates $x_{i}$ for $i=1,...,N$ to select the next set of evaluations. The iterative procedure ends when some stopping criterion is met: computational or experimental resources are expended, a maximum number of iterations is reached, or a satisfactory value of $f$ is achieved.
The imperfect surrogate model $\hat{f}$ provides a probabilistic prediction $y$ of the objective function value(s) for one or more candidates. This distribution over possible values of $y$ for candidate $x$, denoted $p(y|x)$, may be described by a machine learning model such as a Gaussian process or Bayesian neural network. We may alternately have a surrogate model that does not form a continuous probability distribution but still enables sampling, e.g., through deep ensembling 29 or Monte Carlo dropout 30. Without loss of generality, we consider the prediction to be an integral over possible model parameters $\bm{\theta}$:
@@
where $p(\bm{\theta})$ does not truly have to represent a prior, but can represent a posterior distribution given some training data. We consider all discrete candidates $x_{i}\in\mathcal{X}$ to be deterministic and fixed.
Figure 1: Exploitative batch design by maximizing the likelihood of including the optimum, in the context of a single Bayesian optimization iteration. First, a probabilistic surrogate model is trained on acquired data. Second, samples are obtained from the joint posterior distribution over all candidates. When direct posterior sampling is impossible or inefficient, a multivariate Gaussian may be modeled from the true posterior to enable approximate posterior sampling. Third, we estimate from these samples the probability that each candidate is the true optimum. Fourth, the batch is populated with candidates most likely to be optimal; in doing so, the proposed strategy maximizes the probability that the batch contains the true optimum. In addition to the sampling-based approach visualized here, we describe alternative methods to approximate acquisition scores in Section 2.3.2.
@@ 2.2 Deriving an acquisition function for optimal batch design
Most acquisition strategies in BO aim to balance exploitation and exploration. Exploitation prioritizes selections that are most likely to achieve the highest oracle score, while exploration is intended to prevent a search from getting “stuck” in local optima. Conceptually, exploration is expected to contribute to a more reliable surrogate model and thereby benefit the optimization in the long run; a classic failure mode of BO within continuous design spaces is the oversampling of a single mode of the surrogate model posterior 31. In contrast, exploitation selects the best candidates at a given iteration without consideration of the impact on future iterations. We pursue a batch acquisition strategy that is motivated by exploitation, optimizing expected performance in the immediate iteration as if the optimization could be stopped at any time. This is a realistic setting for molecular discovery campaigns; when the cost of evaluating $f$ varies across compounds in the design space, the number of iterations or oracle budget may be uncertain when the optimization begins. Our acquisition strategy is visualized within the context of one BO iteration in Figure 1.
@@
We focus on continuous surrogate models with low but non-zero observation noise (e.g., Gaussian processes), such that the probability of two inputs having the exact same output is 0. Therefore, we may assume that the events $\{\hat{f}(x^{*})=\hat{f}(x_{i})\}_{x_{i}\in\mathcal{X}}$ are mutually exclusive. In our estimation of Eq. 3, we will treat $\{\hat{f}(x^{*})=\hat{f}(x_{i})\}$ as equivalent to $\{x^{*}=x_{i}\}$, leading to the presumed mutual exclusivity of the events $\{x^{*}=x_{i}\}_{x_{i}\in\mathcal{X}}$ and the following result:
$\Pr(x^{*}\in\mathcal{X}_{acq})$ $=\sum_{x_{i}\in\mathcal{X}_{acq}}\Pr(x^{*}=x_{i})$ (4)
$=\sum_{x_{i}\in\mathcal{X}_{acq}}P(f(x^{*})=f(x_{i}))$ (5)
$=\sum_{x_{i}\in\mathcal{X}_{acq}}P(f(x_{i})>f(x_{j})~{}\forall~{}j\neq i).$ (6)
We elaborate on scenarios where a unique optimum cannot be assumed in Section S1. An appropriate strategy to optimize the objective in Eq. 4 is to approximate $\Pr(x^{*}=x_{i})$ for each candidate $x_{i}$ and select the top $b$ candidates based on the resulting approximations. We refer to this acquisition strategy as qPO (multipoint Probability of Optimality), which is deceptively simple to apply in practice. Consider a candidate $x_{i}$ in a list of candidates $\mathbf{x}$—designs that have not been previously acquired. Using the corresponding predictions $\mathbf{y}$ modeled by the surrogate posterior, we may estimate each $\Pr(x^{*}=x_{i})$ as the following integral or expectation:
@@
$=\mathbb{E}_{\mathbf{y}\sim p(\mathbf{y}|\mathbf{x})}\left[\mathbbm{1}_{y_{i}=\text{max}(\mathbf{y})}\right],$ (8)
where each $\alpha_{i}$ represents an acquisition score for $x_{i}$. A special outcome of this batch design strategy is that qPO is naturally a sum over individual acquisition scores. Batch acquisition functions that are extensions of sequential ones, such as multipoint expected improvement (qEI) 21, typically require evaluation of the entire batch at once. Myopic construction of large batches can only approximately optimize these batch-level acquisition functions, risking the selection of suboptimal batches. In contrast, qPO can be optimized without the issue of non-additivity and resulting approximations faced by many myopic batch construction strategies.
@@ 2.3.1 Monte Carlo integration to approximate acquisition scores
Similar to the Monte Carlo approximations of alternative batch acquisition functions 32, the expectation in Eq. 8 may be approximated with $M$ samples from the surrogate model posterior:
$\alpha_{i}$ $\approx\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}_{y_{i}^{(m)}=\text{max}(\mathbf{y}^{(m)})},$ (9)
where $\mathbf{y}^{(m)}=[y_{1}^{(m)},y_{2}^{(m)},...,y_{N}^{(m)}]$ is the $m$th sample from the surrogate model posterior for $N$ candidates. In practice, these may be samples from the posterior of a Gaussian process, predictions based on sampled parameters $\bm{\theta}^{(m)}$ from a Bayesian neural network, or predictions based on Monte Carlo dropout. The result in Section S2 shows that candidates with a high probability of optimality are very likely to appear in the batch even with modest $M$, making a Monte Carlo estimate reasonable. Algorithm 1 summarizes the implementation of qPO using Monte Carlo sampling.
@@
     Compute joint posterior $p(\mathbf{y}|\mathbf{x},\bm{\theta})$ over unacquired candidates $\bf{x}\in\mathcal{X}$
@@
     Evaluate $f(x_{i})$ for all $x_{i}\in\mathcal{X}_{acq}^{*}$triangleright Call the oracle
@@
The only requirement for this approximation of acquisition scores is that the posterior can be sampled. However, efficient and widespread algorithms exist to sample from Gaussian distributions 33, 34. The computational cost of posterior sampling may be reduced by sampling from a multivariate Gaussian that approximates the true posterior. This approximate posterior may be obtained from an arbitrary number of approximate samples from the true posterior. For example, one might obtain a multivariate Gaussian posterior based on the predictions of a neural network ensemble; this could apply for an ensemble of any model architecture and for an arbitrary number of models in the ensemble. Specifically, given a smaller number ($<M$) of samples, one can construct an empirical mean and covariance matrix from which additional samples can be drawn.
@@
If very few candidates are perceived by the model as having a substantial probability of being optimal, $\Pr(x^{*}=x_{i})$ may be estimated to be zero for many candidates for finite $M$. Filling a batch of size $b$ may emerge as a challenge if there are fewer than $b$ non-zero acquisition scores. A simple way to address this issue is to fill the remainder of the batch using an alternative metric (e.g., greedy or upper confidence bound). Methods designed for rare event estimation 35, 36, 37 may alternatively be implemented to assign acquisition scores to candidates with small probabilities of optimality.